OpenStudy (anonymous):

Can someone solve this for me? (3/5 * -1/5 + 4/5 * 2/5)^3

OpenStudy (anonymous):

the answer is 11/125 but not sure how they got it.

OpenStudy (anonymous):

(3/5 * -1/5) = -3/25 (4/5 * 2/5) = 8/25 -3/25 + 8/25 = 5/25 = 1/5 1/5^3 = 1/125 Don't see how it becomes 11 on the numerator. It is impossible as nothing cubed is 11.

OpenStudy (anonymous):

sorry you are right, my key was sticky! yes it is 1/125